# CriticBench: Evaluating Large Language Models as Critic

### v3

@@ 5.2 Overall Analysis of LLMs
As shown in Table 1 and Table 2, GPT-4 significantly outperforms other LLMs on most critique dimensions, approaching the reference critiques (a score of $8$). Besides, open-source LLMs are approaching state-of-the-art closed-source LLMs. For example, InternLM2-20B-Chat surpasses GPT-3.5-turbo on overall subjective scores ($6.20>5.89$). Surprisingly, InternLM2-7B-Chat achieves comparable performance ($5.66$) with much bigger LLMs on the subjective evaluation, like Qwen-72B-Chat and DeepSeek-67B-Chat, but losing to them on the objective evaluation. Critique-tuned LLMs outperform Llama-2 series models while still lagging behind some powerful open-source LLMs, like the InternLM2 series. Besides, reward models like UltraRM-13B achieve impressive performance in scoring the quality of responses, significantly outperforming GPT-3.5-turbo and Claude models.
5.3 Relationship with Tasks
Effective critiques usually require domain knowledge and understanding of given tasks. Thus, we also analyze the relationship between critique ability and tasks in Table 3.
Feedback
LLMs achieve much higher scores in the first five tasks than on math reasoning and coding tasks (average subjective scores $4.92>3.19$ and average objective scores $27.29>18.53$), indicating that math and code tasks are more challenging for critiques. Responses in CodeExec are slightly easier to critique than those in CodeNE, implying a positive effect of code execution information for self-improvement on coding tasks.
Comparison-based Feedback
Similarly, the first five tasks are much easier to critique than mathematics and coding tasks. Besides, LLMs achieve the best scores on QA, chat, and summary tasks, while average scores on coding tasks are the worst.
Correction
Mathematics tasks are more challenging on objective evaluation than coding tasks, and CodeExec are more accessible to revise than those in CodeNE due to richer information in executions. The translation is the most challenging task because professional domain knowledge is required, while harmlessness is the easiest to refine since most LLMs have been trained to avoid harmful generations Bai et al. (2022b); Lee et al. (2023).
Meta-Feedback
LLMs achieve much higher consistency with human judgments on code and mathematics tasks than other $5$ tasks (average score $23.06>3.12$). LLMs perform much worse on translation and summary tasks. This observation proves that automatically evaluating critiques on mathematics and code tasks is more reliable. Moreover, as shown in Table 2, although GPT-4 achieves a very high correlation with human judgments ($62.90$) on the meta-feedback dimension, its gap to the average human level ($79.03$) is $16.13$, indicating an apparent divergence between the GPT-4 and human judgments in assessing critique quality, which has also been observed but not quantitatively verified in recent works Wang et al. (2023c).
Tasks $\boldsymbol{F_{s}(I,R)}$ $\boldsymbol{F_{c}(I,R_{a},R_{b})}$ $\boldsymbol{CR}$ $\boldsymbol{F_{s}(F_{s})}$
@@ 5.3 Relationship with Tasks
Table 3: Average performance of all LLMs on $9$ tasks. Two Avg. rows represent the average scores of the first $5$ tasks and the last $4$ tasks, respectively.
5.4 Difficulty of Critique Dimensions
As shown in last rows of Table 1 and Table 2, the comparison-based feedback presents the highest level of difficulty ($4.58$), followed by the feedback dimension (average score $4.89$), proving that comparison mostly requires accurate analysis of both responses, which is more complex than the single feedback dimension. As a higher-level critique dimension, meta-feedback is more challenging than the feedback dimension ($22.97<35.75$).
Moreover, the correction is the easiest critique dimension with our annotated feedback (average score $7.12$). To understand how feedback affects the quality of refinements, we replace the annotated feedback with Empty feedback (EAvg. column) and LLM’s Self-feedback (SAvg. column) for the correction dimension, and several conclusions can be summarized: (1) most LLMs achieve the best performance with our annotated feedback in CriticBench, proving the effectiveness of our human annotations; (2) high-quality feedback is beneficial to refinement on subjective evaluation ($\textrm{Avg}(7.12)>\textrm{SAvg}(5.48)>\textrm{EAvg}(5.16)$), while low-quality feedback might negatively affect refinement on objective evaluation ($\textrm{SAvg}(13.01)<\textrm{EAvg}(14.44)<\textrm{Avg}(43.66)$). This observation proves that LLMs struggle in self-improvement on challenging mathematics and coding tasks , aligning with recent findings Huang et al. (2024); (3) since GPT-4’s average score on subjective evaluation of the feedback dimension is slightly better than annotated feedback $8.09>8$ (first row in Table 13), it is reasonable that GPT-4 achieves best performance with its feedback. On the contrary, GPT-4’s average feedback score on mathematics and coding tasks is $7.53<8$ (first row in Table 13), leading to worse performance on the objective evaluation ($52.17<69.67$).
@@ 5.5 Relationship with Response Quality
Effective self-improvement needs to identify which kinds of responses are hard to critique. Thus, we analyze the relationship between response quality and critique ability given the Table 1 and Table 2. Most LLMs perform worst on high-quality responses for the feedback dimension, indicating that subtle flaws in these responses pose a significant challenge. In contrast, correct and medium-quality responses are the easiest to critique on the subjective and objective evaluation, respectively. Besides, GPT-4’s critique ability on correct responses $6.07$ is much worse than other responses. This observation also quantitatively confirms recent findings that GPT-4 might modify better answers worse Stechly et al. (2023); Zhang et al. (2024); Huang et al. (2024). For the correction critique dimension, most LLMs perform worst on medium-quality responses while achieving the best performance on high-quality and low-quality responses. For the comp-feedback critique dimension, it is reasonable that samples with distinct qualities are much easier for critics than the harder samples, i.e., two responses with close quality. For the meta-feedback dimension, most LLMs achieve higher consistency with human judgments on high- and low-quality responses than medium-quality responses [^3] Responses here denote the generations $R$ to be critiqued, not the generated critiques $F_{s}$ in the meta-feedback dimension., indicating the difficulty of measuring critiques on medium-quality responses.
5.6 Relationship with Scaling Law
Recent works Wei et al. (2022) have shown that as the scale of models gradually increases, their capability across various tasks also improves. It is also a crucial question whether critique ability obeys this scaling law. As shown in Tables 1 and 2, most LLMs’ critique ability (objective and subjective evaluation) also improves steadily as the number of parameters increases, as observed in the Qwen series, Llama-2 series, DeepSeek series, Mistral series, and InternLM2 series LLMs. For example, as the parameters of Qwen series LLMs increase from 7B to 72B, the overall subjective scores exhibit a progressive increase from $4.63$ to $6.01$, and overall objective scores increase from $34.87$ to $58.48$.
6 Related Work
Self-improvement by Critique Ability
Critique ability has been widely used for LLM’s self-improvement in the following two stages: (1) Inference stage: through prompting LLMs to generate feedback and refinements sequentially, LLMs could iteratively improve their generations Pan et al. (2023); Madaan et al. (2023); Zhang et al. (2024); Saunders et al. (2022); Yao et al. (2023); Fernandes et al. (2023); Huang et al. (2024); Saunders et al. (2022); (2) Training stage: the refinements could be used for the supervised fine-tuning (SFT) or preference learning (RLHF) to improve LLM. They are also widely used in the RLAIF pipeline Bowman et al. (2022); Saunders et al. (2022); Yuan et al. (2024); Bai et al. (2022b); Xu et al. (2023). Meanwhile, minor human principles are still beneficial to LLM self-alignment. Sun et al. (2023, 2024b).
Self-improvement highly relies on LLM’s critique ability. However, there are few studies for measuring critique ability comprehensively, especially the correction and meta-feedback dimensions, impeding the in-depth understanding of the bottleneck of self-improvement. This paper proposes the CriticBench to fill this gap.
Improvement of Critique Ability
Open-source LLMs are proven to have poor critique ability Ye et al. (2023); Wang et al. (2023c). Recently, numerous works have been proposed to improve Llama-2 models by fine-tuning on critiques generated by GPT-4 Li et al. (2024); Cui et al. (2023); Kim et al. (2024); Ke et al. (2023). However, their evaluation is limited in the task scenarios, critique dimensions, critique formats, and test data size.
Application of Critique Ability
So far, critique ability is widely used for automatic evaluation of response quality, frequently utilizing GPT-4 OpenAI (2023); Liu et al. (2023c); Fu et al. (2023) or critique-tuned LLMs Ke et al. (2023); Li et al. (2024), whereas they are proved unstable for some complex task scenarios Wang et al. (2023c); Zhang et al. (2024). Besides, evaluating open-source LLMs’ critique ability is overlooked in previous works Li et al. (2024); Cui et al. (2023). Our work not only proposes a systematic benchmark but also comprehensively evaluates closed-source and open-source LLMs on critique ability Touvron et al. (2023); DeepSeek-AI (2024); Xu et al. (2024); Chiang et al. (2023); Team (2023); Bai et al. (2023), showing several intriguing phenomena.
@@ 7 Conclusion and Future Work
In this paper, we introduce a systematic and comprehensive benchmark, named CriticBench, specifically developed to assess the critique capabilities of LLMs. Furthermore, we conducted an extensive evaluation of both open-source and closed-source LLMs on CriticBench. The in-depth analysis shed light on the critique capabilities of these models concerning critique dimensions, tasks, response qualities, and model sizes, significantly promoting the development of LLM’s self-improvement and scalable oversight. Such an analysis also pinpoints potential pitfalls in self-improvement and scalable oversight of LLMs, especially when the LLM becomes more powerful.
@@
In the future, we plan to enhance our benchmark and evaluation protocol in several key areas: (1) Broadening the scope to include more tasks, like tool-using; (2) Extending the benchmark to encompass other languages, like Chinese; (3) Improving the subjective evaluation protocol to allow for more fine-grained analysis, particularly verifying the claims made in generated feedback.
8 Limitations
@@
CriticBench mainly utilizes the advanced GPT-4 model for subjective evaluation. Despite integrating high-quality annotated reference critiques to guide GPT-4 toward more accurate assessments, it’s essential to acknowledge that the model’s evaluations may not always align perfectly with human judgment. For example, although GPT-4 achieves strong consistency with average human judgments of $62.90$ (Table 2), the gap between GPT-4 and the average human level is $16.13$, showing an apparent divergence. While LLMs have yet to reach the level of precision of human annotation - they currently represent the most effective approach for balancing the trade-offs between evaluation cost and quality. It is still a significant challenge to accurately and automatically evaluate critiques across all scenarios. Recognizing this, we aim to address these issues in our future work by progressively refining our benchmark and evaluation protocols.
9 Ethical Considerations
@@
Most task inputs in CriticBench are collected from publicly available datasets, free from any possible harm toward individuals or groups. Moreover, humans carefully select and process the responses and critiques generated by LLMs to secure privacy and confidentiality. No personal identification information is involved. However, it should be noted that the task input, responses, and critiques in the Anthropic-HHH dataset Bai et al. (2022a) of the harmlessness task contain harmful materials and hate speech. Despite the risks involved, it is essential to disclose this research fully, and materials in the Anthropic-HHH dataset have been widely used for safety research in the LLM community.
Appendix A Source Data for Different Tasks
The details of selected datasets for $9$ tasks are listed in Table 4, covering the well-known NLP tasks (translation, summary, and question answering), reasoning tasks (mathematics and coding), and alignment (general chat and harmlessness). These datasets’ test sets are used for CriticBench construction, avoiding data contamination.
Tasks Source From Test Data Num. Translation WMT20 MLQE Specia et al. (2020) $100$ Chat ChatArena Alpaca-Eval Li et al. (2023) $50$ each QA OBQA Mihaylov et al. (2018) CommonQATalmor et al. (2019) PIQA Bisk et al. (2020) $35$ each Harmlessness HHH Bai et al. (2022a) $100$ Summary Summ. HF Stiennon et al. (2020) $100$ Math PoT Math CoT AquA-RAT Ling et al. (2017) MathQA Amini et al. (2019) GSM8K Cobbe et al. (2021) NumGLUE Mishra et al. (2022) TheoremQA Chen et al. (2023b) $20$ each Code w/. exec Code w/o. exec MBPP Austin et al. (2021) HumanEval Chen et al. (2021) $50$ each
Table 4: Source of $9$ tasks in CriticBench. Most tasks contain diverse samples from multiple test sets.
Appendix B List of Used LLMs
Table 5: The list of used LLMs for generating responses and critiques.
Appendix C Likert Score for Responses
Figure 4: Human annotated Likert scores (1-7) reflecting response quality across tasks. Since execution leaks the information on response quality, correct responses are not collected for the CodeExec task.
Appendix D Statics of CriticBench
The comparison between CriticBench and existing benchmarks can be found in Table 6, which proves the advantages of our proposed CriticBench over previous benchmarks for critique evaluation. The number of samples in CriticBench are shown in Table 7. Compared with existing benchmarks for critique evaluation, our proposed CriticBench contains $5,972$ samples, leading to a more stable and reliable assessment. Specifically, CriticBench includes $3,608$ natural language critique samples (test $,01$ 010 and dev $2598$) and $2,892$ scalar-based critique samples (test $1,214$ and dev $1,688$), which human experts well annotate. The scale of CriticBench is much bigger than existing benchmarks for measuring LLM’s critique ability, such as Auto-J Li et al. (2024), UltraFeedback Cui et al. (2023), and Shepherd Wang et al. (2023c).

### v4

@@ 6.4 Overall Analysis of LLMs
Table 6: Critique-tuned LLMs results in feedback dimension.
Models Sub. Obj
Llama-2-13B 3.70 30.61
Llama-2-70B 4.12 32.79
Auto-J-13B 4.21 36.05
UltraCM-13B 4.12 21.51
TigerScore-13B 3.31 17.87
As shown in Table 2, GPT-4 significantly outperforms other LLMs on most critique dimensions, while slightly underperforms our human-annotated critiques ($7.81<8$). Surprisingly, open-source LLMs are approaching state-of-the-art closed-source LLMs. For example, InternLM2-20B-Chat surpasses GPT-3.5-turbo on overall subjective scores ($6.20>5.89$). Furthermore, there is a clear relationship that the critique ability of LLMs improves steadily as the number of parameters increases (Table 2), suggesting that the critique ability of LLMs highly correlates with their capability. We also provide a clear diagram to show this relationship in Figure 12 in Appendix P. Beyond the average scores, we also categorize the textual critiques of LLMs into multiple quality intervals for more interpretable analysis, which are described in Appendix M.
Table 7: Reward model objective results in $F_{s}$ and $F_{c}$ dimensions.
Models $\boldsymbol{F_{s}}$ $\boldsymbol{F_{c}}$
GPT-3.5-turbo 51.44 40.67
UltraRM-13B 52.33 54.67
Ziya-7B 25.81 40.00
The results of critique-tuned LLMs in the feedback dimension on the test set are shown in Table 6. It can be found that critique-tuned LLMs fine-tuned from Llama-2-13B significantly outperform even the Llama-2-70B-Chat model, proving the effectiveness of critiques datasets (li2024generative, ; cui2023ultrafeedback, ). The results of representative reward models are shown in Table 7. From these results, it can be found that reward models like UltraRM-13B achieve impressive performance in scoring the quality of responses, significantly outperforming GPT-3.5-turbo. This observation aligns with findings in recent works kim2024prometheus .
Table 8: Two Avg. rows represent the average scores of all LLMs on the first $5$ tasks and the last $4$ tasks.
Tasks $\boldsymbol{F_{s}}$ $\boldsymbol{F_{c}}$ $\boldsymbol{CR}$ $\boldsymbol{F_{s}(F_{s})}$
@@
6.5 Relationship with Task Type
Effective critiques usually require domain knowledge and understanding of given tasks. We analyze the relationship between critique ability and task type in Table 8, which shows the average performance of all evaluated LLMs.
Feedback, Comparison
LLMs achieve much higher scores in the first five tasks than on math reasoning and coding tasks, indicating math reasoning and code tasks are more challenging.
Meta-Feedback
LLMs achieve much higher consistency with human judgments on code and math reasoning tasks, indicating that evaluating textual critiques in math reasoning and code tasks is more reliable.
Table 9: Average performance of evaluated LLMs on test set.
Dimension Sub. Obj.
$\boldsymbol{CR}$ w/ HF 7.12 43.66
$\boldsymbol{CR}$ w/ SF 5.48 13.01
$\boldsymbol{CR}$ w/ EF 5.16 14.44
Correction
Math reasoning tasks are more challenging than coding tasks, and CodeExec is easier to revise than CodeNE due to the richer information in execution results. Except for math reasoning and coding tasks, the translation is the most challenging task because professional domain knowledge is required, while harmlessness is the easiest to refine since most LLMs have been trained to avoid harmful generations (bai2022constitutional, ). Furthermore, we explore the variance in correction quality on reasoning and coding tasks (Obj. in Table 9) and other subjective tasks (Sub. in Table 9). [^2] Subjective tasks represent five tasks (translation, general chat, QA, summary and harmlessness). Specifically, three kinds of feedback are used for correction: (1) Human-annotated Feedback (HF); (2) Empty Feedback (EF), where LLMs are prompted to improve responses without any feedback; and (3) LLMs Self-generated Feedback (SF). As shown in Table 9, it can be found that self-generated feedback is beneficial to corrections on subjective evaluation ($\textrm{HF}(7.12)>\textrm{SF}(5.48)>\textrm{EF}(5.16)$), while it might negatively affect corrections on objective evaluation of math reasoning and coding tasks ($\textrm{SF}(13.01)<\textrm{EF}(14.44)<\textrm{HF}(43.66)$). This observation proves that LLMs struggle in self-improvement on challenging reasoning tasks, aligning with recent findings (huang2024large, ; gou2024critic, ).
@@ 6.6 Relationship with Response Quality
Table 10: Error pattern distribution (%).
Error Pattern Low Med. High
Obvious 74.68 29.48 20.42
Complex 16.46 45.51 31.69
Subtle 8.86 25.00 47.89
Before analyzing the relationship between response qualities and critique ability, it is essential to categorize the error patterns in responses. We highlight that the error patterns are related to the task type, complicating the classification of errors. To conduct a representative analysis of errors in all tasks, human annotators are asked to categorize errors into three patterns, which collectively encompass nearly all the cases: (1) Obvious error is easy to critique and correct, like apparent misuses of words in translation task; (2) Complex error is challenging to correct, regardless of whether critiques are easy to critique, like logical reasoning error in reasoning tasks; (3) Subtle error is hard to critique, while it is usually easier to revise than complex error, like slight misunderstandings of context in general chat. The distribution presented in Table 10 reveals distinct primary errors across different response qualities. More details about these error patterns in each task are described in Appendix N.
Table 11: Average performance of LLMs on the different response qualities (test set).
Quality Subjective Objective
$\boldsymbol{F_{s}}$ $\boldsymbol{CR}$ $\boldsymbol{F_{s}}$ $\boldsymbol{CR}$ $\boldsymbol{F_{s}(F_{s})}$
Low 5.14 7.17 21.93 46.04 22.73
Medium 4.76 7.08 23.10 40.58 19.78
High 4.66 7.15 20.62 45.19 28.84
Given the distribution of error patterns, we analyze critique ability of LLMs on responses with varying qualities. As shown in Table 11, high-quality responses are the hardest for feedback since they contain lots of subtle errors (Table 10). Note that the medium-quality responses have higher objective feedback scores than low-quality ones, which is inconsistent with our expectations. This phenomenon is because low-quality responses often receive very low human-annotated quality scores (near 1), while the scoring of LLMs tends to be higher, leading to a discrepancy.
For the correction dimension, low-, and high-quality responses are easier to correct than medium-quality due to the most obvious and subtle errors. There are two kinds of qualities for comparison dimension: easy and hard. Most LLMs perform better on easy samples than on hard samples. Specifically, the subjective and objective scores of easy samples are 4.78 and 39.73, respectively, higher than those of hard samples (4.55 and 29.80). For the meta-feedback dimension, LLMs achieve the highest consistency with human judgments on high-quality responses while performing worst on medium-quality responses.
6.7 Relationship with Critique Dimensions
Table 12: Average performance on test set.
Dimen. Sub. Obj.
$\boldsymbol{F_{s}}$ 4.89 35.75
$\boldsymbol{F_{c}}$ 4.58 -
$\boldsymbol{F_{s}(F_{s})}$ 22.97
$\boldsymbol{CR}$ 7.12 -
The average scores of all evaluated LLMs on different critique dimensions are shown in Table 12. Objective scores of comparison and correction are not recorded because they are not correlations. Several conclusions can be made: (1) correction is the easiest critique dimension, followed by feedback, and then comparison. This observation demonstrates that comparison requires accurate analysis of both responses, which is more complex than the feedback dimension; (2) As a high-level critique dimension, meta-feedback is more challenging than the feedback.
@@ 7 Conclusion and Future Work
In this paper, we introduce a comprehensive and reliable benchmark for evaluating the critique abilities of LLMs, named CriticEval. Extensive evaluation first proves the reliability of CriticEval. Then, experimental results reveal the promising potential of open-source LLMs, the effectiveness of critique datasets and intriguing relationships between critique capabilities and crucial factors: task types, response qualities and critique dimensions. These observations significantly promote an in-depth understanding of the critical ability of LLMs and LLM’s self-improvement.
@@
In the future, we plan to enhance our benchmark in several key areas: (1) Broadening the scope to include more tasks, such as tool-using; (2) Extending the benchmark to encompass other languages, like Chinese; (3) Improving the subjective evaluation protocol to allow for more fine-grained analysis; (4) Continue to evaluate LLMs and track their critique ability, like Llama-3 models; (5) Improving the quality of reference critiques by incorporating additional high-quality critiques from advanced LLMs if and only if their quality surpasses the existing reference critiques.
Appendix A Limitations
A.1 Sub-optimal Reference Critiques
Following previous work liu2023alignbench , our work construct CriticEval with the human-in-the-loop annotation pipeline, i.e., multiple human annotators are asked to review and revise the critiques generated by GPT-4 model. Even though we have established a rigorous annotation process to ensure the high quality of annotated critiques, human annotators are inevitably influenced by GPT-4’s initial generated critiques in some more open-ended tasks, like general chat and QA tasks. This problem may result in the quality of revised critiques by human annotation still being sub-optimal. To address this problem, we plan to enhance CriticEval in our next version. Specifically, we will replace existing reference critiques with potential better critiques generated by advanced evaluated LLMs, like Claude-$3$https://www.anthropic.com/news/claude-3-family, if and only if their qualities surpasses reference critiques.
A.2 GPT-4 Model for Subjective Evaluation
@@
CriticEval mainly utilizes the advanced GPT-4-turbo model for subjective evaluation. Despite integrating high-quality annotated reference critiques to guide GPT-4 toward more accurate assessments, it’s essential to acknowledge that the model’s evaluations may not always align perfectly with human judgment. While GPT-4 have yet to reach the level of precision of human annotation - they currently represent the most effective approach for balancing the trade-offs between evaluation cost and quality. It is still a significant challenge to accurately and automatically evaluate critiques across all scenarios. Recognizing this, we aim to address these issues in our future work by progressively refining our benchmark and evaluation protocols.
A.3 More Tasks for Critique
Compared with existing benchmarks cui2023ultrafeedback ; wang2023shepherd ; kim2024prometheus ; sun2024critique , our proposed CriticEval exhibits significant advantages in the diversity of evaluated task scenarios. Although we strive to cover a wide range of diverse generation tasks, there are still some tasks have yet to be considered, such as tool learning (qin2023tool, ), knowledge-intensive tasks (petroni2021kilt, ) and hallucination (zhang2023sirens, ). In our future work, we will continue to include more tasks in the next version of CriticEval.
A.4 More LLMs to be Evaluated
Some newly proposed LLMs and reward models lambert2024rewardbench have not been added yet, like Llama-3444https://ai.meta.com/blog/meta-llama-3/ and Claude-3 series models. Since our conclusions are summarized from evaluation results of over 35 open-source and closed-source LLMs, lacking these LLMs does not affect conclusions. We will continue to evaluate their critique ability in our future work.
Appendix B Ethical Considerations
@@
Most task inputs in CriticEval are collected from publicly available datasets, free from any possible harm toward individuals or groups. Moreover, humans carefully select and process the responses and critiques generated by LLMs to secure privacy and confidentiality. No personal identification information is involved. However, it should be noted that the task input, responses, and critiques in the Anthropic-HHH dataset (bai2022training, ) of the harmlessness task contain harmful materials and hate speech. Despite the risks involved, it is essential to disclose this research fully, and materials in the Anthropic-HHH dataset have been widely used for safety research in the LLM community. All raters have been paid adequate wages. The hourly wage of our human annotators is about 5.69 USD, which is much higher than average hourly wage 3.13 USD on Amazon Mechanical Turk (hara2018data, ).
Appendix C Comparison and Statistics of CriticEval
The comparison between CriticEval and existing benchmarks can be found in Table 13, which proves the advantages of our proposed CriticEval for critique evaluation. Compared with existing benchmarks for critique evaluation, our proposed CriticEval contains $3,608$ textual natural language critique samples (textual critiques) that are well annotated by multiple human annotators, leading to a more stable and reliable assessment in our subjective evaluation. The scale of objective evaluation data (scalar data) in our dataset is second only to Chat Arena [^5] https://hf-mirror.com/datasets/lmsys/lmsys-arena-human-preference-55k and SummEval (fabbri2021summeval, ). However, compared to these two datasets, our dataset contains a more diverse range of critique dimensions and tasks (nine diverse tasks). Moreover, the statistics of CriticEval in the test and dev set are shown in Table 14.
Table 13: Statistics of existing critique benchmarks, meta-evaluation benchmarks (scalar-valued critique evaluation), and CriticEval. NL and Scalar denote natural language feedback and scalar-valued feedback, i.e., the preference label or Likert score (pan2023automatically, ). CriticBench (luo2023critique, ; lin2024criticbench, ) contain two kinds of response quality (correct and wrong). The responses in some benchmarks are not unclassified, and we set them as unclassified (-). Scalar-valued critiques in Auto-J li2024generative are from its Eval-P, and textual critiques are from Eval-C split.